Of the 600 tickets sold to a movie, 520 of them were full price tickets costing $8 each. If the gate receipts were $4,320 what did a student ticket cost
Found 2 solutions by zhagi17, macston:
Answer by zhagi17(16)
Answer by macston(5194)

.
Gate from full price tickets=(Full Price)(No. of tickets)
Gate from full price tickets=($8)(520)=$4160
.
Student receipts=(Total receipts)-(full price receipts)
Student receipts=$4320-$4160=$160
.
Number of student tickets=(Total tickets)-(No. of full price tickets)
Number of student tickets=600-520=80
.
Student ticket price=(Student receipts)/(No. of student tickets)
Student ticket price=$160/80=$2
.
ANSWER: A student ticket cost $2.